# Invisible Servoing: a Visual Servoing Approach with Return-Conditioned Latent Diffusion

### v1

@@ Abstract
In this paper, we present a novel visual servoing (VS) approach based on latent Denoising Diffusion Probabilistic Models (DDPMs). Opposite to classical VS methods, the proposed approach allows reaching the desired target view, even when the target is initially not visible. This is possible thanks to the learning of a latent representation that the DDPM uses for planning and a dataset of trajectories encompassing target-invisible initial views. The latent representation is learned using a Cross-Modal Variational Autoencoder, and used to estimate the return for conditioning the trajectory generation of the DDPM. Given the current image, the DDPM generates trajectories in the latent space driving the robotic platform to the desired visual target. The approach is applicable to any velocity-based controlled platform. We test our method with simulated and real-world experiments using generic multi-rotor Uncrewed Aerial Vehicles (UAVs). A video of our experiments can be found at https://youtu.be/yu-aTxqceOA.
@@ 1 Introduction
Figure 1: Real world experiment of an invisible servoing maneuver performed by a multi-rotor UAV; top: camera view.
@@
  • validation of the invisibility servoing in simulation and real-world using generic multi-rotor UAVs (Fig. 1).
The approach can generalize to different robotic platforms. Indeed, despite the simulated and real UAV having different dynamics, the velocity actions planned by the DDPM are tackled by the specific motion controller.
@@ 2 Related Work
An alternative to classical VS is Direct VS (DVS). DVS approaches operate directly on the visual input [4, 5], eliminating the need for feature extraction and tracking. The visual input includes information like pixel intensity values or image gradients, which are used to determine the robot’s control commands. By operating on the whole image, DVS formulates control laws based on the similarity between two images. However, the optimization problem at the core is highly nonlinear and large in dimension, therefore the convergence basin of DVS approaches is small, i.e., such methods cannot be used to regain visibility far from the target image. Our approach works directly in the image space but can converge from any view experienced in the dataset, even if the target is not visible.
Recently, a new approach towards offline Deep Reinforcement Learning (DRL) in robotics has emerged exploiting DDPMs for planning and control [16, 10]. By conditioning the generation process on suitable expected returns, the DDPM can strategically generate task-specific trajectories that maximize cumulative rewards over a prediction horizon. This has been demonstrated to even achieve better results than traditional offline DRL methods specially in tasks that require a high degree of flexibility and dexterity [17]. To the best of our knowledge, we are the first to address VS using DDPMs in the latent space, leveraging the recent Cross-Modal Variational Autoencoders (VAEs) [8, 9] to learn a representation that allows defining a suitable return for the VS task.
@@ 3.1 Denoising Diffusion Probabilistic Models
Denoising diffusion probabilistic models (DDPMs) [18, 19] are a class of probabilistic generative models, such as generative adversarial networks [20], variational autoencoders [21], and score-matching models [22, 23]. DDPMs are parametrized Markov chains that learn data distributions out of which new data can be generated from random noise by performing a step-wise denoising of the random vectors. The denoising process, often referred to as the reverse process, is learned from the dataset with the goal of reverting the forward process of gradually adding noise to the original data. The forward process is shown in (1):
@@ 4 Methodology
However, the high dimensionality of image data makes the methods introduced in [16, 10] not directly applicable due to the need for generating extremely high-dimensional state trajectories, which is computationally untractable. Our novel framework builds upon the observation that the input data, i.e., the camera images, contains redundant information and that the relevant one lie in a low-dimensional space. Therefore, we can simplify the problem of VS with DDPM by taking inspiration from the state-of-the-art latent DDPM [28]. The problem of VS can then be solved in two steps: (i) learn a low-dimensional representation of the input data using a Cross-Modal Variational Autoencoder, and (ii) learn to generate latent space trajectories using a DDPM (see Fig. 3).
@@ 4.1 Dimensionality Reduction using Cross-Modal Variational Autoencoder
To address the challenge of high dimensionality in image data and learn a compact and regularized latent space a Cross-modal Variational Autoencoder (CM-VAE) [8] is utilized. This approach has been shown to reduce the data dimensionality effectively, while ensuring the generated latent space is smooth, continuous, consistent, and robust against variations in visual information [8, 9]. The idea behind the CM-VAE is to integrate multiple data modalities into a single latent space. The CM-VAE processes two primary modalities: (i) RGB image data $\mathbf{x}_{\text{rgb}}$ and (ii) the visual feature’s relative pose $\mathbf{x}_{\text{f}}$ with respect to the robot’s coordinate frame, i.e., $\mathbf{x}_{\text{f}}=[r,\theta,\phi,\gamma]$, expressed using the spherical coordinates and the yaw.
The input image $\mathbf{x}_{\text{rgb}}$ is projected by a deep neural network-based encoder $q^{\text{rgb}}_{\boldsymbol{\theta}_{E}}$ of parameters $\boldsymbol{\theta}$ into a normal distribution $\mathcal{N}(\mu,\sigma^{2})$, from which the latent vector $\mathbf{z}$ is sampled [^3] Similarly to classical VAEs, the encoder learns the mean and covariance of the Gaussian distribution over the latent variables. To sample from such a distribution a latent vector $\mathbf{z}$ the reparametrization trick is used [21].. The latent vector $\mathbf{z}$ is used to reconstruct each data modality dedicated deep neural network-based decoders $p^{\text{rgb}}_{\boldsymbol{\theta}_{D}}$ for images and $p^{\text{f}}_{\boldsymbol{\theta}_{d}}$ for feature poses. More specifically, the full latent vector is used in the decoding of the image data, while for decoding the feature relative pose only the first four scalars of the latent vector are used, i.e., $\tilde{\mathbf{z}}=[z_{1},z_{2},z_{3},z_{4}]$ is decoded to $r,\theta,\phi$ and $\gamma$ respectively. For the encoder and the two decoders we utilize the same neural network architecture introduced in [8].
@@ 4.2 Latent DDPMs for Visual Servoing
As shown in Fig. 3, the image encoder $q_{\text{rgb}}$, trained to minimize the CM-VAE loss function in (9), is used to encode the image data into the regularized latent space for the diffusion model to operate in. In particular, the DDPM task is now to learn to generate latent state trajectories of the type:
@@
To train the DDPM model, we project the image trajectories into the latent space using the trained CM-VAE to obtain the trajectories in the latent space. For conditional sampling, the diffusion model is conditioned on a desired outcome or quality expressed as a desired return to achieve. Therefore, the dataset is augmented with the corresponding returns achieved by each trajectory. Then, the latent DDPM is trained by a straightforward modification of classifier-free guidance objective introduced in (6) jointly with the latent inverse model (see (8)):
@@ 4.2.1 Return Estimation
In our VS task, the diffusion model is trained on a dataset in which the reward at any step in the trajectory is equal to minus the distance between the visual feature and the robot. Therefore, during planning, if the distance between the visual feature and the robot is known, the return can be estimated by assuming a constant-velocity motion of the robot from the start to the target image. The position vector between the visual feature and the robot at the start and end of the trajectory can be calculated using the feature decoder $p^{\text{f}}_{\boldsymbol{\theta}_{d}}$ of the CM-VAE, which takes the encoded image latent vector z as input and outputs the vector $[r,\theta,\phi,\gamma]$. The spherical coordinates $[r,\theta,\phi]$ can be converted to Cartesian coordinates to get the position vectors $\vec{r}_{0}$ and $\vec{r}_{N}$ at the start and end of the trajectory respectively, where $N$ is the length of the prediction horizon. The distance vector $\vec{d}$ between the start $\vec{r}_{0}$ and target position $\vec{r}_{N}$ can be calculated as $\vec{d}=\vec{r}_{0}-\vec{r}_{N}$. Assuming that the robot moves with a constant velocity of magnitude $v$ along this vector $\vec{d}$ from the start position to the target position, the total travel time can be calculated as $T=||\vec{d}||/v$. Therefore, the time between each step $k$ in the trajectory can be calculated as $\Delta t=T/N$. The position vector $\vec{r}_{k}$ between the visual feature and the robot at each step $k$ in the trajectory can be computed by sliding along the vector $\vec{d}$ with a step size of $v\Delta t$ starting from the start position $\vec{r}_{0}$:
$\vec{r}_{k}$ $=\left(\vec{d}-vk\Delta t\hat{d}\right)+\vec{r}_{N}$ (13)
$=\left(||\vec{d}||\hat{d}-vk\Delta t\hat{d}\right)+\vec{r}_{N}$ (14)
$=\left(||\vec{d}||-vk\Delta t\right)\hat{d}+\vec{r}_{N}\,,$ (15)
@@
where $\hat{d}$ is the unit vector in the direction of $\vec{d}$. The return is then estimated by summing the discounted rewards along the trajectory where the reward at each step $k$ is the negative of the distance $||\vec{r}_{k}||$ between the visual feature and the robot at that step:
$\text{R}_{est}=-\sum_{k=0}^{N}\gamma^{k}||\vec{r}_{k}||.$ (16)
@@
The constant velocity magnitude $v$ in the return estimation algorithm affects the calculated return. Therefore, the suitable value of $v$ should be chosen based on the robot’s dynamics and the environment. To estimate this velocity, the dataset is used to calculate this velocity with two different approaches. The first approach is to calculate the average velocity of the robot in the dataset and use this velocity for any planning regardless of the start and target images. The second approach is to predict the velocity based on the linear and yaw angular distances between the start and target images where, again, these are calculated using the feature decoder of the CM-VAE. This velocity prediction is obtained from a linear regression model (linear 3D plane) that has been fit to the dataset. The model takes the linear and yaw angular distances as input and outputs the predicted velocity. Both approaches to select the velocity magnitude $v$ yield very similar results.
5 Simulations and Experimental Validation
The main focus of this work is to propose a novel VS approach based on latent DDPMs, that leverages a cross-modal regularized latent representation to estimate effective return values for conditional sampling. For illustration purposes, a color map of a generated latent trajectory is shown in Fig. 4.
@@ 5 Simulations and Experimental Validation
To test our approach, we have considered two experimental scenarios: a simplified simulated environment of an empty square room (Fig. 5(a)), and the Aerial Robotics Lab at our department (Fig. 5(b)). In the former, the camera resolution is $128\times 128$ pixels and the FoV $1.21\,\mathrm{rad}$. The dataset consists of quadrotor trajectories, counting up to 700k image-datapoint pairs and containing: captured images, the visual feature pose w.r.t the robot, the reward at each time step, the robot’s linear and angular velocities used as high-level control inputs. The visual target is the Aruco tag visible in Fig. 5(a). In the simulated Aerial Robotics Lab, the camera resolution is $320\times 180$ pixels, and the FoV $1.21\,\mathrm{rad}$. These values correspond to one of the streaming configurations available for the Intel Realsense D435i camera, which we used for real-world experiments. The dataset consists of hexarotor and quadrotor trajectories, counting up to 700k image-datapoint pairs. The visual target is the RAM logo attached to the wall, visible in Fig. 5(b). In the datasets, the return values are normalized between -1.0 and 1.0.
Our experiments are designed to show the capabilities of our method and support our claim that it can effectively reach a desired target view, even if this is not visible at the start. In particular, the experiments focus on (i) the performance of the return estimation heuristic built upon the latent representation given by the CM-VAE, (ii) the performance of the diffusion-based controller working in a receding horizon fashion, (iii) the transfering of the control stragies from the simulation to the real-world without any need for real-world data during training.
@@ 5.1 Return estimation
The first experiment is designed to show the effectiveness of the return estimation module, which highly affects the performance of the latent DDPM. Planning takes place in the latent space, the trajectory of images is then decoded to visualize and interpret the planned trajectory in the image space.
@@
5.2 Receding horizon control
The next set of experiments is designed to support the claim that our approach can be applied in a receding horizon fashion, allowing to reach the visual target in the challenging scenario where the target is not initially visible.
Fig. 7 shows the closed-loop experiment for the hexarotor in the lab environment. The DDPM samples a trajectory, then the first five control actions are executed before sampling a new trajectory again. In sequence, the estimated return values are 0.46, 0.52, 0.60, 0.64, 0.68, 0.73, consistently growing as the UAV approaches the target. As it is well known, the closed-loop experiment allows the correction of the uncertainties in the model, thus improving the result in terms of the error at the final pose.

### v2

@@ Abstract
In this paper, we present a novel visual servoing (VS) approach based on latent Denoising Diffusion Probabilistic Models (DDPMs), that explores the application of generative models for vision-based navigation of UAVs (Uncrewed Aerial Vehicles). Opposite to classical VS methods, the proposed approach allows reaching the desired target view, even when the target is initially not visible. This is possible thanks to the learning of a latent representation that the DDPM uses for planning and a dataset of trajectories encompassing target-invisible initial views. A compact representation is learned from raw images using a Cross-Modal Variational Autoencoder. Given the current image, the DDPM generates trajectories in the latent space driving the robotic platform to the desired visual target. The approach has been validated in simulation using two generic multi-rotor UAVs (a quadrotor and a hexarotor). The results show that we can successfully reach the visual target, even if not visible in the initial view.
@@ 1 Introduction
(a) Gazebo simulation environment, camera view, and highlight of the visual target.
(b) Five receding horizon steps of the returned-conditioned latent diffusion allow the UAV to reach the visual target.
Figure 1: An overview of the maneuver generated by the proposed approach to steer the UAV to the desired visual target, using the images of the onboard camera.
@@
  • validation of the invisibility servoing in simulation using generic multi-rotor UAVs (Fig. 1).
The approach can generalize to different robotic platforms, as the velocity actions planned by the DDPM are tackled by the specific motion controller. We have tested our approach in simulation, applying the planned actions in a receding horizon fashion to allow for online correction of model uncertaintes. Future work includes experiments in real-world and more complex scenarios, and investigation of an end-to-end pipeline.
@@ 2 Related Work
An alternative to classical VS is Direct VS (DVS). DVS approaches operate directly on the visual input [4, 5], eliminating the need for feature extraction and tracking. The visual input includes information like pixel intensity values or image gradients, which are used to determine the robot’s control commands. By operating on the whole image, DVS formulates control laws based on the similarity between two images. Albeit novel approaches have been proposed to overcome the difficulty of a large-scale highly nonlinear problem, e.g. leveraging Discrete Orthogonal Moments [16], those approaches still require the visual target to be visible.
Recently, a new approach towards offline Deep Reinforcement Learning (DRL) in robotics has emerged exploiting DDPMs for planning and control [17, 10]. By conditioning the generation process on suitable expected returns, the DDPM can strategically generate task-specific trajectories that maximize cumulative rewards over a prediction horizon. This has been demonstrated to even achieve better results than traditional offline DRL methods specially in tasks that require a high degree of flexibility and dexterity [18]. Leveraging the ideas of planning in the image space [19] and learning a latent representation of the images [20], this paper explores the use of DDPMs in the latent space for DVS. This has been possible building on the recent Cross-Modal Variational Autoencoders (VAEs) [8, 9] and the return estimation algorithm we implemented.
@@ 3.1 Denoising Diffusion Probabilistic Models
Denoising diffusion probabilistic models (DDPMs) [21, 22] are a class of probabilistic generative models, such as generative adversarial networks [23], variational autoencoders [24], and score-matching models [25, 26]. DDPMs are parametrized Markov chains that learn data distributions out of which new data can be generated from random noise by performing a step-wise denoising of the random vectors. The denoising process, often referred to as the reverse process, is learned from the dataset with the goal of reverting the forward process, which gradually adds noise to the original data $\mathbf{x}_{t}$ until convergence to a Gaussian distribution $\mathcal{N}(\mathbf{x}_{t};\boldsymbol{\mu},\mathbf{\Sigma})$ with zero mean $\boldsymbol{\mu}$ and unit variance $\mathbf{\Sigma}$. The forward process is detailed as follows:
@@ 4 Methodology
However, the high dimensionality of image data makes the methods introduced in [17, 10] not directly applicable due to the need for generating extremely high-dimensional state trajectories, which is computationally untractable. Our novel framework builds upon the observation that the input data, i.e., the camera images, contains redundant information and that the relevant one lies in a low-dimensional space. Therefore, we can simplify the problem of VS with DDPM by taking inspiration from the state-of-the-art latent DDPM [31]. The problem of VS can then be solved in two steps: (i) learn a low-dimensional representation of the input data using a Cross-Modal Variational Autoencoder, and (ii) learn to generate latent space trajectories using a DDPM (see Fig. 3).
@@ 4.1 Dimensionality Reduction using Cross-Modal Variational Autoencoder
To address the challenge of high dimensionality in image data and learn a compact and regularized latent space, a Cross-modal Variational Autoencoder (CM-VAE) [8] is utilized. This approach has been shown to reduce the data dimensionality effectively, while ensuring the generated latent space is smooth, continuous, consistent, and robust against variations in visual information [8, 9]. The idea behind the CM-VAE is to integrate multiple data modalities into a single latent space. The CM-VAE processes two primary modalities: (i) RGB image data $\mathbf{x}_{\text{rgb}}$ and (ii) the visual feature’s relative pose $\mathbf{x}_{\text{f}}$ with respect to the robot’s coordinate frame, i.e., $\mathbf{x}_{\text{f}}=[r,\theta,\phi,\gamma]$, expressed using the spherical coordinates and the yaw.
The input image $\mathbf{x}_{\text{rgb}}$ is projected by a deep neural network-based encoder $q^{\text{rgb}}_{\boldsymbol{\theta}_{E}}$ of parameters $\boldsymbol{\theta}$ into a normal distribution $\mathcal{N}(\boldsymbol{\theta};\boldsymbol{\mu},\boldsymbol{\Sigma})$, from which the latent vector $\mathbf{z}$ is sampled [^3] Similarly to classical VAEs, the encoder learns the mean and covariance of the Gaussian distribution over the latent variables. To sample from such a distribution a latent vector $\mathbf{z}$ the reparametrization trick is used [24].. The latent vector $\mathbf{z}$ is used to reconstruct each data modality by two dedicated deep neural network-based decoders $p^{\text{rgb}}_{\boldsymbol{\theta}_{D}}$ for images and $p^{\text{f}}_{\boldsymbol{\theta}_{d}}$ for feature poses. More specifically, the full latent vector is used in the decoding of the image data, while for decoding the feature relative pose only the first four scalars of the latent vector are used, i.e., $\tilde{\mathbf{z}}=[z_{1},z_{2},z_{3},z_{4}]$ is decoded to $r,\theta,\phi$ and $\gamma$ respectively. For the encoder and the two decoders we utilize the same neural network architecture introduced in [8].
@@ 4.2 Latent DDPMs for Visual Servoing
As shown in Fig. 3, the image encoder $q^{\text{rgb}}_{\boldsymbol{\theta}_{E}}$, trained to minimize the CM-VAE loss function in (9), is used to encode the image data into the regularized latent space for the diffusion model to operate in. In particular, the DDPM task is now to learn to generate latent state trajectories of the type:
@@
To train the DDPM model, we project the image trajectories into the latent space using the trained CM-VAE to obtain the trajectories in the latent space. For conditional sampling, the diffusion model is conditioned on a desired outcome or quality expressed as a desired return to achieve. To make this possible, the dataset is augmented with the corresponding returns achieved by each trajectory. Then, the latent DDPM is trained by a straightforward modification of classifier-free guidance objective introduced in (6), jointly with the latent inverse model (see (8)):
@@
Figure 4: Linear regression model prediction of the velocity.
@@ 4.2.1 Return Estimation
In our VS task, the diffusion model is trained on a dataset in which the reward at any step in the trajectory is equal to minus the distance between the visual feature and the robot. Therefore, during planning, if the distance between the visual feature and the robot is known, the return can be estimated by assuming a constant-velocity motion of the robot from the start to the target image. The position vector between the visual feature and the robot at the start and end of the trajectory can be calculated using the feature decoder $p^{\text{f}}_{\boldsymbol{\theta}_{d}}$ of the CM-VAE, which takes the encoded image latent vector z as input and outputs the vector $[r,\theta,\phi,\gamma]$. The spherical coordinates $[r,\theta,\phi]$ can be converted to Cartesian coordinates to get the position vectors $\mathbf{r}_{0}$ and $\mathbf{r}_{N}$ at the start and end of the trajectory respectively, where $N$ is the length of the prediction horizon. The distance vector $\mathbf{d}$ between the start $\mathbf{r}_{0}$ and target position $\mathbf{r}_{N}$ can be calculated as $\mathbf{d}=\mathbf{r}_{0}-\mathbf{r}_{N}$. Assuming that the robot moves with a constant velocity of magnitude $v$ along this vector $\mathbf{d}$ from the start position to the target position, the total travel time can be computedDo we w as $T=||\mathbf{d}||/v$. Therefore, the time between each step $k$ in the trajectory can be calculated as $\Delta t=T/N$. The position vector $\mathbf{r}_{k}$ between the visual feature and the robot at each step $k$ in the trajectory can be computed by sliding along the vector $\mathbf{d}$ with a step size of $v\Delta t$ starting from the start position $\mathbf{r}_{0}$:
$\mathbf{r}_{k}$ $=\left(\mathbf{d}-vk\Delta t\mathbf{\hat{d}}\right)+\mathbf{r}_{N}$ (13)
$=\left(||\mathbf{d}||-vk\Delta t\right)\mathbf{\hat{d}}+\mathbf{r}_{N}\,,$ (14)
@@
where $\mathbf{\hat{d}}$ is the unit vector in the direction of $\mathbf{d}$. The return is then estimated by summing the discounted rewards along the trajectory where the reward at each step $k$ is the negative of the distance $||\mathbf{r}_{k}||$ between the visual feature and the robot at that step:
$\text{R}_{est}=-\sum_{k=0}^{N}\gamma^{k}||\mathbf{r}_{k}||.$ (15)
@@
The constant velocity magnitude $v$ in the return estimation algorithm affects the calculated return. Therefore, the suitable value of $v$ should be chosen based on the robot’s dynamics and the environment. To estimate this velocity, the dataset is used to calculate this velocity with two different approaches. The first approach is to calculate the average velocity of the robot in the dataset and use this velocity for any planning regardless of the start and target images. The second approach is to predict the velocity based on the linear and yaw angular distances between the start and target images where, again, these are calculated using the feature decoder of the CM-VAE. This velocity prediction is obtained from a linear regression model (linear 3D plane) that has been fit to the dataset as shown in Fig. 4. The model takes the linear and yaw angular distances as input and outputs the predicted velocity. The two approaches to select the velocity magnitude $v$ yield very similar results.
@@
5 Simulations
The main focus of this work is to propose a novel VS approach based on latent DDPMs, that leverages a cross-modal regularized latent representation to estimate effective return values for conditional sampling. In particular, we want to test the capability of the DDPM-based controller to recover visibility and approach a given target view. Therefore, the simulation environments encompass a visual target, that the robot has to reach being spawned at random positions in the rooms.
To test our approach, we have considered two experimental scenarios: a simplified simulated environment of an empty square room (Fig. 5(a)), and the Aerial Robotics Lab at our department (Fig. 5(b)). In the former, the camera resolution is $128\times 128$ pixels and the FoV $1.21\,\mathrm{rad}$. The dataset consists of quadrotor trajectories, counting up to 700k image-datapoint pairs and containing: captured images, the visual feature pose w.r.t the robot, the reward at each time step, the robot’s linear and angular velocities used as high-level control inputs. The visual target is the Aruco tag visible in Fig. 5(a). In the simulated Aerial Robotics Lab, the camera resolution is $320\times 180$ pixels, and the FoV $1.21\,\mathrm{rad}$. The dataset consists of hexarotor and quadrotor trajectories, counting up to 700k image-datapoint pairs. The visual target is the RAM logo attached to the wall, visible in Fig. 5(b). In the datasets, the return values are normalized between -1.0 and 1.0.
The Gazebo simulations show the capabilities of our method to reach a desired target view, even if this is not visible at the start. The evaluation focuses on (i) the performance of the return estimation heuristic built upon the latent representation given by the CM-VAE, (ii) the performance of the diffusion-based controller working in a receding horizon fashion.
@@ 5.1 Return estimation
As the dynamics are not modeled as hard constraints by the proposed approach, the effectiveness of the return estimation module is essential to generate dynamically feasible trajectories. Once a suitable return value has been calculated, planning takes place in the latent space. For illustration purposes, a color map of a generated latent trajectory is shown in Fig. 6. Once a latent space trajectory has been generated similar to the one in Fig. 6, the decoder $p^{\text{rgb}}_{\boldsymbol{\theta}_{D}}$ is used to visualize the trajectory in the image space converting each latent vector to an image.
Figure 8: 100 samples of generated UAV trajectories in open-loop fashion, along with their mean and standard deviation. The visual target is not visible from the initial location (-1.0, 0.5). The latent DDPM is capable of recovering the view of the target.
@@
To give an intuition of the probabilistic nature of the proposed approach, a batch of 100 open-loop trajectories generated by the DDPM is shown in Fig. 8, along with the highlight of the initial target-blind view and of a sample successful final view. We analyzed the residual error between the final pose and the pose where the desired target view was captured. Fig. 9 shows the distribution of the residual error in the form of histograms.
5.2 DDPM-based controller
Scenario Return
1 2 3 4 5 6
Box 0.62 0.69 0.71 0.76 0.87 0.88
Lab 0.46 0.52 0.60 0.64 0.68 0.73
Table 1: Estimated return values at different consecutive samplings of the receding horizon control in the two simulated scenarios.
Figure 9: Error between the final pose and the pose where the desired target image was captured in open-loop simulations: (A) position; (B) yaw.
In the following, we analyze the performance of the proposed approach, showing that the receding horizon paradigm is the key to reaching the visual target and that the return estimation algorithm consistently estimates lower returns the more the robot approaches the goal.
In Fig. 1, we have shown the closed-loop experiment for the hexarotor in a simulated lab environment. The DDPM samples a trajectory, then the first five control actions are executed before sampling a new trajectory again. As shown in Tab. 1, the estimated return is consistently growing as the UAV approaches the target, both in the lab and in the box simulation environments. As it is well known, the closed-loop experiment allows the correction of the uncertainties in the model. As a consequence, the average residual error can be drastically reduced: on one of the initial positions, it improved from 1.22 m to 0.10 m in position, and from 0.43 rad to 0.40 rad in yaw.